OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (anonymous):

Replace the x in function f with the expression for g. Then simplify.

OpenStudy (anonymous):

does anybody know how to do it ^^

OpenStudy (anonymous):

okay thank you

OpenStudy (lieutenantgeneral):

f(x) x = 2/9 g(x) x = 3

OpenStudy (anonymous):

number three is the correct answer I need to find the formula to the problem Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). and I have no clue how does anyone ? 1. -9x - 2 2. -9x + 5 3. -9x + 25 4. -9x + 27

OpenStudy (anonymous):

anyone ^^ ??

OpenStudy (anonymous):

I really need help on this

OpenStudy (anonymous):

Replace x in f(x) with g(x) F(g(x))=9(-x+3)-2=-9x+27-2=-9x+25

OpenStudy (anonymous):

thank you sooooo very much !!!!!!!!

OpenStudy (anonymous):

U r welcome

OpenStudy (anonymous):

can anyone help me with, Let f(x) = 2x + 2. Solve f-1(x) when x = 4. the correct answer is 1 I just need to find out the formula to find 1. does anyone have a solution ?

OpenStudy (anonymous):

I really need help on this ?? ^^

OpenStudy (anonymous):

R su about the answer?re

OpenStudy (anonymous):

Yes the choices are 1 3 4 10 1 Is correct

OpenStudy (anonymous):

I need the formula

OpenStudy (anonymous):

Oh need to find the inverse of the function

OpenStudy (anonymous):

how do I do that ?

OpenStudy (anonymous):

This is my last problem I have no clue how to do it can you help me ?

OpenStudy (anonymous):

X=2y+2 2y=x-2 F-1(x)=x-2/2 f-1(4)=4-2/2=2/2=1

OpenStudy (anonymous):

Thank you very much !!!!!!

OpenStudy (anonymous):

U r welcome:)

OpenStudy (anonymous):

Let f(x) = x2 - 81. Find f-1(x). does anyone know the answer to this equation ??

OpenStudy (anonymous):

f-1(x)=sqrt(x+81)

OpenStudy (anonymous):

thank you !!

OpenStudy (anonymous):

U r welcome :) medal?? If u don't mind :D

OpenStudy (anonymous):

how do I give you a medal ?? And what is the slope of the line between (3, -4) and (-2, 1)? I need help with that problem ??

OpenStudy (anonymous):

@31356

OpenStudy (31356):

Correct! :D